OpenStudy (anonymous):

Given sec(theta) = -4/3 and 90 degrees < (theta) <180 degrees ; find sin2(theta)

OpenStudy (usukidoll):

90 degrees < (theta) <180 degrees tells us that we are restricted in the 2nd quadrant... so all sine values are going to be positive...

OpenStudy (anonymous):

If sec(theta) = -4/3 cos(theta) = -3/4 Use sin^2(x)+cos^2(x)=1 to solve for sin^2(x). Plug in -3/4 for cos(x), square it, and subtract the result from 1.

OpenStudy (usukidoll):

\[\sin^2x\] or \[\sin2x\] for the last part of the question. They have different meanings.

OpenStudy (usukidoll):

\[\sin^2\theta............ \sin2\theta\]

OpenStudy (anonymous):

its Sin2theta

OpenStudy (usukidoll):

alright... sin2theta has an identity \[2\sin\theta \cos\theta\]

OpenStudy (usukidoll):

given that sec(theta) = -4/3 sec(theta) =\[\frac{1}{-\frac{4}{3}} \rightarrow \frac{-3}{4}\]

OpenStudy (usukidoll):

it's negative due to the fact that we have the restriction of 90 degrees <theta<180 degrees which means 2nd quadrant and only sine values are positive now let's ignore the - for a bit and find out what our sine really is so cosine is adjacent/hypotenuse so that means cosine 3/4

OpenStudy (usukidoll):

 so we need to use the pythagorean theorem a^2+b^2=c^2

OpenStudy (usukidoll):

in this case we solve for a. a^2+3^2=4^2 a^2+9=16 a^2 = 16-9 a^2 = 7 a = \[\sqrt{7}\] ugh one problem... I know we can't pick the negative result for a because negative sides are nonexistent and doesn't make sense... I don't think decimals can be used either... not sure.

OpenStudy (usukidoll):

sec(theta) = -4/3 cos(theta) = -3/4 in 2nd quadrant hmmm... cos(theta) 3/4 = adjacent/hypotenuse and then solve for a... that should be ...why the heck do I have square root of 7 ? sine theta is \[\frac{\sqrt{7}}{4}\] ... I thought all sides are supposed to have whole numbers

OpenStudy (usukidoll):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

It's perfectly possible to have a side length that isn't a whole number example: